The armature of a small generator consists of a flat, square coil with 120 turns and sides
with a length of 1.6 cm. The coil rotates in a magnetic field of 0.075T. What is the
angular speed of the coil if the maximum emf produced is 24.0mV?
1
Expert's answer
2020-04-08T10:26:52-0400

Flux is given by "\phi =BA\cos\omega t" "instantaneous\ emf = e(t) = - N d(\phi)/dt = - Nw (BA) [- sin wt]"

maximum emf= "NwBA"

"0.024=120\times w\times0.075\times0.016^2\\w=10.41rad/sec"